OpenStudy (anonymous):

help please. Is perpendicular to y = -2/3x + 4 and passes through the point (2, -3)

OpenStudy (anonymous):

First find slope of given line comparing it with y = mx + c Where m = slope and c = y intercept There is relation between slopes of perpendicular line as m1 * m2 = -1 Where m1= slope of given line m2= slope of second line From this you would get slope m2 Now we are having slope and point so use slope - point form of line as y - y1 = m ( x - x1 ) where m = slope and ( x1 , y1 ) is given point Plug all values into above equation and try to simplify to get equation of desired line .......